There are various methods in literature that countries use in calculating national income. Assume a simple economy that produces only two goods; maize and fish. Using 2015 as the base year.

Goods 2015 2016 2017 2018
Maize Quantities 700 900 1000 950
Fish Quantities 500 1000 900 750
Price of Maize K65 K70 K80 K90
Price of Fish K50 K70 K90 K120
Nominal GDP
Real GDP


a) Fill in the missing cells in the above table. Clearly show your work.

b) Find the growth rate of nominal GDP for 2016, 2017 and 2018.
c) Find the growth rate of real GDP (using 2015 K) for 2016, 2017 and 2018.

d) Compute inflation using GDP deflator (using 2015 K) for 2016, 2017, and 2018.
e) What are the major differences between the GDP deflator and the CPI?
1
Expert's answer
2018-08-15T08:58:08-0400
a) Fill in the missing cells in the above table.
Nom GDP 70500 133000 161000 175500
Real GDP 70500 108500 110000 99250
b) The growth rate of nominal GDP is:
2016: (133000/70500 - 1)*100% = 88.65%,
2017: (161000/133000 - 1)*100% = 21.05%,
2018: (175500/16100 - 1)*100% = 9%,
c) The growth rate of real GDP (using 2015 K) is:
2016: (108500/70500 - 1)*100% = 53.9%,
2017: (110000/108500 - 1)*100% = 1.38%,
2018: (99250/110000 - 1)*100% = -9.02%,
d) Inflation using GDP deflator (using 2015 K) is:
2016: (133000/108500 - 1)*100% = 22.58%,
2017: (161000/110000 - 1)*100% = 46.36%
2018: ( 175500/99250 - 1)*100% = 76.83%
e) The major differences between the GDP deflator and the CPI are that the GDP deflator compares the price level in the current year to level in the base year, but CPI indicates the prices of a representative basket of commodities procured by the consumers.